OpenStudy (anonymous):

The table below shows two equations: Equation 1 |2x − 3| + 5 = 4 Equation 2 |5x + 3| − 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 2, −3.2. The solutions to equation 1 are x = 1, 2 and equation 2 has no solution. The solutions to equation 1 are x = 1, 2 and equation 2 has solutions x = 2, −3.2.

OpenStudy (campbell_st):

what do you know about absolute value...?

OpenStudy (anonymous):

It is the number's distance from zero.

OpenStudy (campbell_st):

as an example.... can you have \[\left| x \right| = -10\]

OpenStudy (campbell_st):

that's correct... its distance from zero

OpenStudy (anonymous):

But I have no idea how to solve the problems, so I can't answer the question. :c

OpenStudy (campbell_st):

ok... so look at the 1st equation... \[\left| 2x-3 \right|+5 = 4\] the 1st step is to subtract 5 from both sides of the equation.

OpenStudy (campbell_st):

so you get \[\left| 2x -3 \right| = -1\] this does not have a solution as you can have an absolute value equal to a zero.

OpenStudy (anonymous):

So |2x−3|= -1?

OpenStudy (campbell_st):

equation 2 add 10 to both sides \[\left| 5x + 3 \right| = 13\] this makes sense... so you now can proceed so you need o solve for the positive and negative 5x + 3 = 13 and 5x + 3 = -13 hope it helps

OpenStudy (anonymous):

Yes, thank you, I think I got it now.